# Gauging Non-Invertible Symmetries: Topological Interfaces and Generalized Orbifold Groupoid in 2d QFT

### 22 Dec 2023 v2

@@ 3.5 Module Categories and Orbifold Groupoid for ${\rm\bf Rep}(D_{8})$
From the complete list of haploid algebras $(A,m)$ in ${\rm\bf Rep}(D_{8})$ found in the last subsection, we can immediately read off the corresponding NIM-irreps from Section 3.1 and they are listed in Table 2. Below, we will explain how we obtain the rest of the gauging data summarized in Table 2 as well as the generalized orbifold groupoid for ${\rm\bf Rep}(D_{8})$. We will be brief about similar steps that we have gone over in more detail for ${\rm\bf Rep}(H_{8})$ in Section 3.3 and focus on the new features for ${\rm\bf Rep}(D_{8})$. We first note that the F-symbols for ${\rm\bf Rep}(D_{8})$ from (3.5) respects the $S_{3}$ automorphism of the fusion ring [90],
@@
We can then go down the list of $L$ in Table 2 to determine the dual fusion category ${}_{A}{\rm\bf Rep}(D_{8})_{A}$ for each algebra $A$ and the corresponding $L$, using well-known results in gauging non-anomalous abelian subgroups of finite groups [9, 105, 10, 33]. When $L=D_{8}$ (with or without discrete torsion), we get back ${\rm\bf Rep}(D_{8})$ as is evident from the above discussion. All other subgroups of $D_{8}$ are abelian. When $L$ is a non-normal subgroup (i.e. $L=\langle s\rangle\,,\langle sr\rangle$), this discrete gauging again creates the non-invertible symmetry category ${\rm\bf Rep}(D_{8})$, where the non-invertible duality TDL comes from the the direct sum of TDLs in the nontrivial $L$-orbit. When $L$ is an abelian normal subgroup, then the dual symmetry is again a group of order 8 containing normal subgroup $\hat{L}\cong L$. The full group structure depends on whether $L$ gives a nontrivial extension of the quotient $D_{8}/L$ by $L$ [^28] A group extension given by the short exact sequence $1\to N\to G\to Q\to 1\,,$ (3.34) is nontrivial if the extension class $H^{2}(Q,N)$ is taken to be nontrivial. Equivalently, a trivial group extension corresponds to a split exact sequence which means $G=N\rtimes Q$ is a semi-direct product (which includes the direct product $G=N\times Q$ as a special case). Here for $D_{8}$ the only nontrivial extension arises for the exact sequence (3.26), in which case we obtain the well-known result that the dual symmetry is ${\rm\bf Vec}_{{\mathbb{Z}}_{2}^{3}}^{\alpha}$ where the nontrivial extension leads to an anomaly characterized by the following 3-cocycle in $H^{3}({\mathbb{Z}}_{2}^{3},U(1))={\mathbb{Z}}_{2}^{7}$ [107], [^29] In the original context [107], this explains the equivalence between the twisted $2+1d$ abelian Dijkgraaf-Witten gauge theory ${\rm DW}[{\mathbb{Z}}_{2}^{3}]_{\alpha}$ (where ${\alpha}$ specifies the TQFT action and the untwisted non-abelian gauging theory ${\rm DW}[D_{8}]$.
@@
The full (connected) groupoid ${\rm\bf BrPic}$ is presented in Figure 12. It consists of three nodes representing the Morita equivalent fusion categories ${\rm\bf Rep}(D_{8})$, ${\rm\bf Vec}_{D_{8}}$, and ${\rm\bf Vec}_{{\mathbb{Z}}_{2}^{3}}^{\alpha}$ connected by Morita equivalence relations (invertible bimodule categories). There are a total of $21624\times 3^{2}=$ 216 invertible bimodule categories.
@@ 4 Examples of Generalized Gauging in CFT
[…]
@@ 4.1.1 Operator Content and Verlinde TDLs in ${\rm Ising}^{2}$ CFT
[…]
@@
[…]
@@
[…]
@@ 4.1.2 Infinite Generalized Symmetries on $c=1$ Orbifold Branch
[…]
@@
and identification (orbifold) by the ${\mathbb{Z}}_{2}^{C}$ symmetry that sends $\vartheta$ to $-\vartheta$. The radius $R$ in (4.5) parametrize the continuous orbifold branch of the conformal moduli space at $c=1$ [109]. T-duality leads to an equivalence relation among the orbifold CFTs by
@@
where $s,r$ generate the $D_{8}$ invertible symmetry (see (3.7)) on the orbifold branch that comes from the nontrivial extension of the commutant ${\mathbb{Z}}_{2}^{\vartheta}\times{\mathbb{Z}}_{2}^{\varphi}\subset G_{\rm circ}$ of ${\mathbb{Z}}_{2}^{C}$ by the dual (magnetic) ^$\hat{{\mathbb{Z}}}_{2}$ symmetry as a consequence of the nontrivial mixed anomaly (as in (3.35)) for ${\mathbb{Z}}_{2}^{\vartheta}\times{\mathbb{Z}}_{2}^{\varphi}\times{\mathbb{Z}}_{2}^{C}$ [33]. Consequently the dual symmetry to ${\mathcal{C}}_{\rm circ}$ given by the symmetry group (4.7) on the orbifold branch is,
@@
[…]
@@
[…]
@@
[…]
$A=\mathds{1}\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}_{{\rm\bf Rep}({D_{8}})}\,,\quad(\mathds{1}\oplus\eta\oplus\eta^{\prime}\oplus\eta\eta^{\prime}\oplus 2{\mathcal{V}}_{{\rm\bf Rep}({D_{8}})})_{1,2}\,,\quad$ (4.19)
@@
${\mathcal{L}}_{{\pi\over 4},0}^{2}=\mathds{1}\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}_{{\rm\bf Rep}({D_{8}})}\,,\quad$ (4.20)
@@
[…]
@@ 4.2 Generalized Gauging and Orbifold Groupoid in Irrational CFT
and thus is equivalent to gauging the ${\mathbb{Z}}_{2}\subset D_{8}$ subgroup generated by $\eta=sr$. The rest of the gaugings in Table 1 follow from sequential gauging with invertible symmetries. Similarly, the top maximal algebra of ${\rm\bf Rep}(D_{8})$ in Table 2 satisfies, [^36] We have chosen a particular $S_{3}$ triality frame for this to hold as algebra objects, where the RHS carries the canonical algebra structure for objects of the form ¯$\overline{\mathcal{L}}\otimes A\otimes{\mathcal{L}}$.
@@ 4.3.2 Gauging Binary Algebras in $SU(2)_{10}$ CFT
For illustration, let us consider gauging binary algebras in the Wess-Zumino-Witten (WZW) CFT for $SU(2)$ symmetry. As an RCFT, it is constructed from $k+1$ unitary integrable representations of the affine Kac-Moody algebra $\widehat{\mathfrak{su}(2)}_{k}$ for $k\in{\mathbb{Z}}_{\geq 1}$, labeled by $0\leq j\leq{k\over 2}$ (see [65] for details). The RCFT is completely specified by a choice of modular invariant; this modular invariant then falls into the famous ADE classification, which includes the $A_{n}$-type or diagonal modular invariant at all $k=n-1$, the $D_{n}$-type or charge-conjugacy modular invariant at even $k=2n-4$ for $n\geq 4$, and the $E_{6},E_{7},E_{8}$type exceptional modular invariants at $k=10,16,28$, respectively [113, 114]. In particular, the diagonal modular invariant RCFT is particular simple and commonly known as the $SU(2)_{k}$ WZW CFT. Its operator spectrum consists of $k+1$ scalar primary operators in one-to-one correspondence with the unitary integrable representations of $\widehat{\mathfrak{su}(2)}_{k}$. Correspondingly there are $k+1$ Verlinde TDLs that we label as ${\mathcal{L}}_{j}$, all of which are self-dual, and the Frobenius-Schur indicators are given by $\epsilon_{j}=(-1)^{2j}$. The fusion rules of the Verlinde TDLs are
@@
and we also know that ${\mathcal{D}}_{1}^{\prime\prime}={\mathcal{D}}_{1}{\mathcal{D}}_{1}^{\prime}$ has the trivial FS indicator since ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{1}^{\prime}$ are related by parity-flip and must have identical FS indicator. The multiplication morphism for the dual binary algebra (4.48) is then given by
$m_{\mathds{1}\mathds{1}}^{\mathds{1}}=m_{\mathds{1}{\mathcal{D}}_{1}^{\prime\prime}}^{{\mathcal{D}}_{1}^{\prime\prime}}=m_{{\mathcal{D}}_{1}^{\prime\prime}\mathds{1}}^{{\mathcal{D}}_{1}^{\prime\prime}}=m_{{\mathcal{D}}_{1}^{\prime\prime}{\mathcal{D}}_{1}^{\prime\prime}}^{\mathds{1}}=\frac{1}{\sqrt{3+\sqrt{3}}}\,,~{}m_{{\mathcal{D}}_{1}^{\prime\prime}{\mathcal{D}}_{1}^{\prime\prime}}^{{\mathcal{D}}_{1}^{\prime\prime}}=\sqrt{\sqrt{3}-1}\,.$ (4.51)

### 25 Mar 2026 v3

@@ 3.5 Module Categories and Orbifold Groupoid for ${\rm\bf Rep}(D_{8})$
From the complete list of haploid algebras $(A,m)$ in ${\rm\bf Rep}(D_{8})$ found in the last subsection, we can immediately read off the corresponding NIM-irreps from Section 3.1 and they are listed in Table 2.282828We thank Yanming Su for pointing out a typo in the table. Below, we will explain how we obtain the rest of the gauging data summarized in Table 2 as well as the generalized orbifold groupoid for ${\rm\bf Rep}(D_{8})$. We will be brief about similar steps that we have gone over in more detail for ${\rm\bf Rep}(H_{8})$ in Section 3.3 and focus on the new features for ${\rm\bf Rep}(D_{8})$. We first note that the F-symbols for ${\rm\bf Rep}(D_{8})$ from (3.5) respects the $S_{3}$ automorphism of the fusion ring [90],
@@
We can then go down the list of $L$ in Table 2 to determine the dual fusion category $_{A}{\rm\bf Rep}(D_{8})_{A}$ for each algebra $A$ and the corresponding $L$, using well-known results in gauging non-anomalous abelian subgroups of finite groups [9, 104, 10, 33]. When $L=D_{8}$ (with or without discrete torsion), we get back ${\rm\bf Rep}(D_{8})$ as is evident from the above discussion. All other subgroups of $D_{8}$ are abelian. When $L$ is a non-normal subgroup (i.e. $L=\langle s\rangle\,,\langle sr\rangle$), this discrete gauging again creates the non-invertible symmetry category ${\rm\bf Rep}(D_{8})$, where the non-invertible duality TDL comes from the direct sum of TDLs in the nontrivial $L$-orbit. When $L$ is an abelian normal subgroup, then the dual symmetry is again a group of order 8 containing normal subgroup $\hat{L}\cong L$. The full group structure depends on whether $L$ gives a nontrivial extension of the quotient $D_{8}/L$ by LL. [^30] A group extension given by the short exact sequence $1\to N\to G\to Q\to 1\,$, (3.34) is nontrivial if the extension class $H^{2}(Q,N)$ is taken to be nontrivial. Equivalently, a trivial group extension corresponds to a split exact sequence which means $G=N\rtimes Q$ is a semi-direct product (which includes the direct product $G=N\times Q$ as a special case). Here for $D_{8}$ the only nontrivial extension arises for the exact sequence (3.26), in which case we obtain the well-known result that the dual symmetry is ${\rm\bf Vec}_{{\mathbb{Z}}_{2}^{3}}^{\alpha}$ where the nontrivial extension leads to an anomaly characterized by the following 3-cocycle in $H^{3}({\mathbb{Z}}_{2}^{3},U(1))={\mathbb{Z}}_{2}^{7}$ [106], [^31] In the original context [106], this explains the equivalence between the twisted $2+1d$ abelian Dijkgraaf-Witten gauge theory ${\rm DW}[{\mathbb{Z}}_{2}^{3}]_{\alpha}$ (where $\alpha$} specifies the TQFT action and the untwisted non-abelian gauging theory ${\rm DW}[D_{8}]$.
@@
The full (connected) groupoid $\rm\bf BrPic$} is presented in Figure 12. It consists of three nodes representing the Morita equivalent fusion categories ${\rm\bf Rep}(D_{8})$, ${\rm\bf Vec}_{D_{8}}$, and ${\rm\bf Vec}_{{\mathbb{Z}}_{2}^{3}}^{\alpha}$ connected by Morita equivalence relations (invertible bimodule categories). There are a total of $24\times 3^{2}=216$ invertible bimodule categories.
@@ 4 Examples of Generalized Gauging in CFT
4.1 Ising2 CFT and Infinite Non-invertible Self-duality
The first example we study is one of the simplest nontrivial CFTs at $d=2$; it is the theory constructed from the tensor product of two copies of the Ising CFT, which we refer to as the Ising2 CFT. As was analyzed in [28, 33], despite its simplicity, this CFT hosts a rich zoo of non-invertible symmetries that provide an ideal playground to study generalized gauging of non-invertible symmetries.
@@ 4.1.1 Operator Content and Verlinde TDLs in ${\rm Ising}^{2}$ CFT
We start by reviewing the operator content of the Ising2 CFT and its generalized symmetries. The Ising2 CFT has conformal central charge $c=1$ and is a rational CFT (RCFT) with respect to the chiral algebra $V_{{\rm Ising}^{2}}\equiv{\rm Vir}_{c={1\over 2}}\times{\rm Vir}_{c={1\over 2}}$. The primary operators with respect to $V_{{\rm Ising}^{2}}$ are given by tensor products of those in each Ising factor,
@@
where $\epsilon_{i}$ is the $h=\bar{h}={1\over 2}$ energy operator and $\sigma_{i}$ is the $h=\bar{h}={1\over 16}$ spin operator. The full operator content (including the descendants) of the Ising2 CFT is captured by its torus partition function
@@
However, unlike a single Ising factor whose symmetry category is finite and fully determined, the Ising2 CFT has many more symmetries due to the much richer operator content; in fact, ${\mathcal{C}}_{{\rm Ising}^{2}}$ contains uncountably infinitely many simple TDLs [28, 31, 33].
@@ 4.1.2 Infinite Generalized Symmetries on $c=1$ Orbifold Branch
The Ising2 CFT comes with an exactly marginal parameter that generates a one-dimensional conformal manifold which corresponds to the orbifold branch of the $c=1$ CFT. Such CFTs are known to be described by a free compact boson $\vartheta\sim\vartheta+2\pi$ with the following action
@@
and identification (orbifold) by the ${\mathbb{Z}}_{2}^{C}$ symmetry that sends $\vartheta$ to $\vartheta$. The radius $R$ in (4.5) parametrize the continuous orbifold branch of the conformal moduli space at $c=1$ [108]. T-duality leads to an equivalence relation among the orbifold CFTs by
@@
where $s,r$ generate the $D_{8}$ invertible symmetry (see (3.7)) on the orbifold branch that comes from the nontrivial extension of the commutant ${\mathbb{Z}}_{2}^{\vartheta}\times{\mathbb{Z}}_{2}^{\varphi}\subset G_{\rm circ}$ of ${\mathbb{Z}}_{2}^{C}$ by the dual (magnetic) $\hat{{\mathbb{Z}}}_{2}$ symmetry as a consequence of the nontrivial mixed anomaly (as in (3.35)) for ${\mathbb{Z}}_{2}^{\vartheta}\times{\mathbb{Z}}_{2}^{\varphi}\times{\mathbb{Z}}_{2}^{C}$ [33]. Consequently the dual symmetry to ${\mathcal{C}}_{\rm circ}$ given by the symmetry group (4.7) on the orbifold branch is,
@@
4.1.3 General Fusion Category Symmetries in Ising2 CFT
Coming back to the Ising2 CFT that sits at the bosonization/fermionization radius $R=2$ on the $c=1$ orbifold branch, we note the symmetry is further enhanced beyond (4.10) by (4.3) [33], and given the quotient of a bi-crossed product (or an exact factorization) category [110, 111],
@@ 4.1.3 General Fusion Category Symmetries in Ising2 CFT
where the second equality follows T-duality on the orbifold branch (which establishes the isomorphism between the Ising2 and its ${\mathbb{Z}}_{2}$ gauging by ${\mathcal{N}}^{2}=\mathds{1}\oplus\eta$).
4.1.4 Generalized Gauging in Ising2 CFT and Infinite Non-invertible Self-duality
Let us now consider generalized gauging in the Ising2 CFT using the symmetries identified in (4.12). For brevity, we mostly focus on two subcategories, namely ${\rm\bf Rep}(H_{8})$ and ${\rm\bf Rep}(D_{8})$ for which we have worked out their universal properties concerning generalized gauging in Section 3, and we will now see how they are realized in a nontrivial physical CFT.
We start with gauging the ${\rm\bf Rep}(H_{8})$ symmetry of Ising2 CFT. The possible distinct gaugings are classified in Table 1. We note immediately that 6 of the 8 algebra objects $(A,m)$ in ${\rm\bf Rep}(H_{8})$ become Morita trivial (with respect to a larger but finite subcategory) in the symmetry category ${\mathcal{C}}_{{\rm Ising}^{2}}$ of the Ising2 CFT, because of the existence of simple TDLs in ${\mathcal{C}}_{{\rm Ising}^{2}}$ that “square” to $A$,
@@ 4.1.4 Generalized Gauging in Ising2 CFT and Infinite Non-invertible Self-duality
Therefore, all except (4.16) gauging of ${\rm\bf Rep}(H_{8})$ in the Ising2 CFT necessarily produce self-duality. [^35] In particular, ${\mathcal{N}}{\mathcal{L}}_{{\pi\over 4},\mp{\pi\over 2}}\cong{\mathcal{N}}{\mathcal{L}}_{{3\pi\over 4},\pm{\pi\over 2}}\,$, (4.18) are the two self-duality TDLs ${\mathcal{D}},{\mathcal{D}}^{\prime}$ identified in the recent work [81] for gauging the unique maximal algebra in a specific ${\rm\bf Rep}(H_{8})$ symmetry of the ${\rm Ising}^{2}$ CFT [33]. They are related by fusing with the invertible generator $s\in D_{8}$ that swaps the two Ising factors and they generate two inequivalent fusion categories as ${\mathbb{Z}}_{2}$ extensions of the ${\rm\bf Rep}(H_{8})$. Such extensions of a fusion category $\mathcal{C}$} by $G$ are classified using the Brauer-Picard groupoid of $\mathcal{C}$} [71]. More explicitly, the inequivalent $G$-extensions correspond to homotopy classes of maps between classifying spaces $BG\to B\underline{\underline{\rm\bf BrPic}}({\mathcal{C}})$ (which consists of a group homomorphism $\rho:G\to{\rm\bf BrPic}({\mathcal{C}})$ together with a suitable 2-cocycle and a 3-cocycle for $G$ that satisfy vanishing obstruction conditions) [71]. Here for $G={\mathbb{Z}}_{2}$ and ${\mathcal{C}}={\rm\bf Rep}(H_{8})$, with the Brauer-Picard groupoid given in Section 3.3 (see around (3.27)), there are 8 inequivalent fusion categories from the ${\mathbb{Z}}_{2}$-extension of ${\mathcal{C}}={\rm\bf Rep}(H_{8})$ parametrized by four distinct choices of $\rho:{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}^{3}$ and the Frobenius-Schur indicator $\epsilon_{\mathcal{D}}$ of $\mathcal{D}$} [81]. In the Ising2 CFT, $\epsilon_{\mathcal{D}}=1$ (see Footnote 33) and two of the remaining four fusion categories (generated by $\mathcal{D}$} and $\mathcal{D}$$^{\prime}$ in [81]) are in one-to-one correspondence with those in (4.18). We thank the authors of [81] for a correspondence that led to a corrected statement above. It is straightforward to check this explicitly at the level of torus partition functions using the explicit form of the multiplication morphism given in Section 3.
One can carry out the parallel analysis for gauging the ${\rm\bf Rep}(D_{8})$ symmetry of the Ising2 CFT, whose algebra objects are listed in Table 2. The 11 Morita equivalence classes of algebra objects for ${\rm\bf Rep}(D_{8})$ descend to 4 Morita equivalence classes with respect to a larger symmetry subcategory. Among them, the Morita trivial class contains, in addition to those in (4.15) and the trivial class in Table 2, the following algebras
$A=\mathds{1}\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}_{{\rm\bf Rep}({D_{8}})}\,,\quad(\mathds{1}\oplus\eta\oplus\eta^{\prime}\oplus\eta\eta^{\prime}\oplus 2{\mathcal{V}}_{{\rm\bf Rep}({D_{8}})})_{1,2}\,$, (4.19)
@@
${\mathcal{L}}_{{\pi\over 4},0}^{2}=\mathds{1}\oplus\eta\eta^{\prime}\oplus{\mathcal{V}}_{{\rm\bf Rep}({D_{8}})}\,$$\quad$ (4.20)
@@
To complete this analysis, let us discuss what happens when we gauge the Morita nontrivial algebras in ${\mathcal{C}}_{{\rm Ising}^{2}}$. Under gauging either algebras in (4.16) (which are Morita equivalent), the Ising2 CFT turns into the compact boson CFT on the circular branch at the same bosonization/fermionization radius $R=2$, which is identified with the Dirac fermion CFT upon bosonization (GSO projection). [^37] The bosonic Dirac CFT is also known as the $U(1)_{4}$ CFT since it is an RCFT with respect to the $U(1)_{4}$ chiral algebra. The dual symmetry for gauging $A=\mathds{1}\oplus\eta\eta^{\prime}$ in ${\rm\bf Rep}(D_{8})$ is ${\rm\bf Vec}_{D_{8}}$ from Table 2, which is an obvious non-anomalous $D_{8}$ subgroup of the invertible symmetry (4.7) on the circular branch generated by the ${\mathbb{Z}}_{4}$ subgroup of $U(1)_{\vartheta}$ together with ${\mathbb{Z}}_{2}^{C}$. On the other hand, gauging $A=(\mathds{1}\oplus\eta\oplus\eta^{\prime}\oplus\eta\eta^{\prime})_{\star}$ in Table 2 generates the anomalous ${\rm\bf Vec}_{{\mathbb{Z}}_{2}^{3}}^{\alpha}$ symmetry that is identified with the ${\mathbb{Z}}_{2}^{\vartheta}\times{\mathbb{Z}}_{2}^{\varphi}\times{\mathbb{Z}}_{2}^{C}$ subgroup of (4.7), which has a mixed anomaly that involves all three ${\mathbb{Z}}_{2}$ factors [33]. Furthermore, gauging the same algebras in ${\rm\bf Rep}(H_{8})$ produces an anomalous dual symmetry ${\rm\bf Vec}_{D_{8}}^{\gamma}$ (see Table 1) that is identified with another $D_{8}$ subgroup of (4.7) generated by ${\mathbb{Z}}_{2}^{C}$ and a ${\mathbb{Z}}_{4}$ translation $\vartheta$ mixed with a ${\mathbb{Z}}_{2}$ translation in $\varphi$.
Finally, under gauging the extra Morita nontrivial class of algebras from ${\rm\bf Rep}(D_{8})$ in (4.22), the Ising2 CFT is mapped to another point on the orbifold branch at $R=4$, which corresponds to the $U(1)_{8}/{\mathbb{Z}}_{2}$ CFT, using the equivalence to the invertible gauging of ${\mathbb{Z}}_{2}^{s}$ [33]. The dual category ${\rm\bf Rep}(D_{8})$ is the familiar symmetry on the entire orbifold branch [33]. When gauging the last Morita nontrivial class in (4.24), we obtain instead the $U(1)_{8}$ CFT on the circular branch at $R=4$, where the dual category ${\rm\bf Vec}_{D_{8}}$ is the same non-anomalous $D_{8}$ symmetry that come from gauging $A=\mathds{1}\oplus\eta\eta^{\prime}$ above (but at a different radius). It is interesting to see how the above discussions are consistent with sequential gauging as there are many possible gauging sequences, especially in the ${\rm\bf Rep}(D_{8})$ case, thanks to the rich groupoid structure.
@@ 4.2 Generalized Gauging and Orbifold Groupoid in Irrational CFT
and thus is equivalent to gauging the ${\mathbb{Z}}_{2}\subset D_{8}$ subgroup generated by $\eta=sr$. The rest of the gaugings in Table 1 follow from sequential gauging with invertible symmetries. Similarly, the top maximal algebra of ${\rm\bf Rep}(D_{8})$ in Table 2 satisfies, [^38] We have chosen a particular $S_{3}$ triality frame for this to hold as algebra objects, where the RHS carries the canonical algebra structure for objects of the form $\overline{{\mathcal{L}}}\otimes A\otimes{\mathcal{L}}$.
@@ 4.3.2 Gauging Binary Algebras in $SU(2)_{10}$ CFT
For illustration, let us consider gauging binary algebras in the Wess-Zumino-Witten (WZW) CFT for $SU(2)$ symmetry. As an RCFT, it is constructed from $k+1$ unitary integrable representations of the affine Kac-Moody algebra $\widehat{\mathfrak{su}(2)}_{k}$ for $k\in{\mathbb{Z}}_{\geq 1}$, labeled by $0\leq j\leq{k\over 2}$ (see [65] for details). The RCFT is completely specified by a choice of modular invariant; this modular invariant then falls into the famous ADE classification, which includes the $A_{n}$-type or diagonal modular invariant at all k=n1k=n-1, the $D_{n}$-type or charge-conjugacy modular invariant at even k=2n4k=2n-4 for $n\geq 4$, and the $E_{6},E_{7},E_{8}$ -type exceptional modular invariants at $k=10,16,28$, respectively [112, 113]. In particular, the diagonal modular invariant RCFT is particular simple and commonly known as the $SU(2)_{k}$ WZW CFT. Its operator spectrum consists of $k+1$ scalar primary operators in one-to-one correspondence with the unitary integrable representations of $\widehat{\mathfrak{su}(2)}_{k}$. Correspondingly there are $k+1$ Verlinde TDLs that we label as ${\mathcal{L}}_{j}$, all of which are self-dual, and the Frobenius-Schur indicators are given by $\epsilon_{j}=(-1)^{2j}$. The fusion rules of the Verlinde TDLs are
@@
and we also know that ${\mathcal{D}}_{1}^{\prime\prime}={\mathcal{D}}_{1}{\mathcal{D}}_{1}^{\prime}$ has the trivial FS indicator since ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{1}^{\prime}$ are related by parity-flip and must have identical FS indicator. The multiplication morphism for the dual binary algebra (4.48) is then given by [^43] We thank Yu Nakayama for pointing out a typo in this formula.
$m_{\mathds{1}\mathds{1}}^{\mathds{1}}=m_{\mathds{1}{\mathcal{D}}_{1}^{\prime\prime}}^{{\mathcal{D}}_{1}^{\prime\prime}}=m_{{\mathcal{D}}_{1}^{\prime\prime}\mathds{1}}^{{\mathcal{D}}_{1}^{\prime\prime}}=m_{{\mathcal{D}}_{1}^{\prime\prime}{\mathcal{D}}_{1}^{\prime\prime}}^{\mathds{1}}=\frac{1}{\sqrt{3+\sqrt{3}}}\,,~m_{{\mathcal{D}}_{1}^{\prime\prime}{\mathcal{D}}_{1}^{\prime\prime}}^{{\mathcal{D}}_{1}^{\prime\prime}}=\sqrt{\frac{\sqrt{3}-1}{3+\sqrt{3}}}\,.$ (4.51)